OpenStudy (ammarah):

Tom, Bob, and Joe are brothers. Their combined ages are 47. Joe was born 3 years before Bob and 4 years after Tom. How old is each brother?

OpenStudy (compassionate):

Tag:

OpenStudy (anonymous):

T + B + J = 47 J = 3+B J = T-4

OpenStudy (ammarah):

what do i do now?

OpenStudy (anonymous):

t = j +4 = 3+b + 4 = b + 7 j = 3 + b b + 7 + b + 3 + b = 47 3b + 10 = 47 3b = 37 b = 12.3 (weird )

OpenStudy (anonymous):

can you confirm that the sum of their ages is 47

OpenStudy (anonymous):

with this : bob is 12.3 joe = 15.3 tom is 19.3 totals 46.9